OpenStudy (anonymous):

1.Write the equation of the line with a slope of -5 and a y-intercept of (0,3).

jimthompson5910 (jim_thompson5910):

y = mx+b y = -5x + 3 Note: the slope is m and the y-intercept is (0, b)

OpenStudy (anonymous):

y=-5x+3?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

yaay thanx

OpenStudy (anonymous):

an y- intercept of (0,3) means "At value x = 0, y must be 3". If you take y = mx + b, where x = 3 and y must be 3, then y = m * 0 + 3 = 3. So like jim says y = -5x + 3

OpenStudy (anonymous):

sorry should be where x = 0